OpenStudy (anonymous):

convert to a mixed number 22/5

OpenStudy (solomonzelman):

\[\huge\color{red}{ \frac{22}{5} } =\huge\color{red}{ \frac{20+2}{5} } =\huge\color{red}{ \frac{20}{5} +} \huge\color{red}{ \frac{2}{5} }=?\]

OpenStudy (solomonzelman):

See how I broke it up?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

So you tell me the answer, what is it?

OpenStudy (anonymous):

4 2/5

OpenStudy (solomonzelman):

YUP!

OpenStudy (anonymous):

actually you could do it pretty easily how many times can 5 fit into 22? 4 so 4 times 5 is 20 and then 2 /5 because you have 2 left over

OpenStudy (anonymous):

so its 4 2/5

OpenStudy (anonymous):

thank you @SolomonZelman

OpenStudy (solomonzelman):

Just another example to show you how this works. \[\huge\color{blue}{ 4 \frac{3}{7} } =\huge\color{blue}{ \frac{(7 \times 4)+3}{7} } \]

OpenStudy (solomonzelman):

You welcome!